OpenStudy (austinl):

Two small, positively charged spheres have a combined charge of 37.0 × 10^-5 C. If each sphere is repelled from the other by an electrostatic force of 0.600 N when the spheres are 2.00 m apart, what is the charge (in Coulombs) on the sphere with the smaller charge?

OpenStudy (austinl):

@broken_symmetry Would you mind taking a looksee at this problem as well? Hate to be a bother.

OpenStudy (austinl):

I have an idea as to how to go about it, but I am not sure that it is possible to solve like that.

OpenStudy (austinl):

q1+q2=37.0E-5C Correct? So q1=37.0E-5C - q2 could you plug in these values and try and solve for them?

OpenStudy (anonymous):

0.6 N * pi * epsilon_0 * 4 m^2 = 6.68e-11 C^2 The product of the charges is 6.68e-11 C^2 The sum of the charges is 37.0e-5 C You now have 2 linearly independent simultaneous equations with two unknowns, so it's solveable.

OpenStudy (anonymous):

Ah, I missed a 4: the product of the charges is 2.67e-10 C^2